SPECTROSCOPY OF THE C^{1}Π- A^{1}Δ TRANSITION OF NH-Ne^{a}

\underline{GALINA KERENSKAYA}, MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The NH-Ne van der Waals complex has been observed in conjunction with the C^{1}Π- A^{1}Δ transition of NH. Low resolution LIF spectra (0.5 cm^{−1} linewidth) exhibit at least three vibronic bands of the complex in the region of the monomer 0- 0 band. The complex bands are spread over a region approximately 30 cm^{−1} wide. The lowest energy band is red-shifted relative to the monomer, indicating that the van der Waals bond strengthens on electronic excitation. Spectra recorded at higher reso lution (0.08 cm^{−1}) show congested rotational structure. Analyses of these data and spectra for ND-Ne will be reported.

aWork supported by National Science Foundation.